# GEM3D: GEnerative Medial Abstractions for 3D Shape Synthesis

### 26 Feb 2024 v1

@@ Abstract.
We introduce GEM3D – a new deep, topology-aware generative model of 3D shapes. The key ingredient of our method is a neural skeleton-based representation encoding information on both shape topology and geometry. Through a denoising diffusion probabilistic model, our method first generates skeleton-based representations following the Medial Axis Transform (MAT), then generates surfaces through a skeleton-driven neural implicit formulation. The neural implicit takes into account the topological and geometric information stored in the generated skeleton representations to yield surfaces that are more topologically and geometrically accurate compared to previous neural field formulations. We discuss applications of our method in shape synthesis and point cloud reconstruction tasks, and evaluate our method both qualitatively and quantitatively. We demonstrate significantly more faithful surface reconstruction and diverse shape generation results compared to the state-of-the-art, also involving challenging scenarios of reconstructing and synthesizing structurally complex, high-genus shape surfaces from Thingi10K and ShapeNet.
@@
Figure 1. We present GEM3D, a neural generative model able to (a) generate skeleton-based representations, or medial abstractions (in red), and 3D surfaces (in green) conforming to the structure and topology information encoded in the skeletons. Our method can be used in various applications, such as (b) reconstructing topologically complex surfaces, (c) synthesizing diverse shape surface samples conforming to a given skeleton, or even (d) synthesizing surfaces from user-specified skeletons far from the ones observed during training.
@@ 1. Introduction
Automated or semi-automated 3D shape synthesis is a significant and challenging problem in geometric modeling, with wide-ranging applications to computer-aided design, fabrication, architecture, art, and entertainment. While early work in this space primarily focused on handwritten models (Prusinkiewicz and Lindenmayer, 1990; Musialski et al., 2013; Merrell et al., 2010), subsequent work employed statistical learning to infer generative design principles from data (Kalogerakis et al., 2012; Fisher et al., 2011). In recent years, a variety of approaches have developed deep neural network-based architectures for 3D synthesis (Xu et al., 2023; Shi et al., 2023; Chaudhuri et al., 2020; Patil et al., 2020). While these methods can capture diverse macro-level appearances, they rarely model shape structure or topology explicitly, relying instead on the representational power of the network to generate plausible-looking voxel grids (Liu et al., 2017), point clouds (Achlioptas et al., 2018a), meshes (Dai and Nießner, 2019), or implicit fields (Chen and Zhang, 2019). Since 3D networks are hampered by the additional resource overheads incurred by the extra dimension compared to 2D image generation networks, they often struggle to model fine detail and connectivity. Some approaches do model part layouts (Li et al., 2017), but are limited in the complexity of the structures they can generate.
At the same time, these prior 3D synthesis methods rarely give artists flexible, precise control. They act more as black boxes for unconditional generation, or reconstruction from images or 3D scans. Recent methods introduce synthesis based on text prompts (Poole et al., 2023; Lin et al., 2023), with remarkable results but only high-level, global control via prompt engineering. 3D character artists have long been accustomed to posing skeleton rigs for accurate character configuration. However, such direct local control and interpretability through intuitive abstractions has had limited success in general 3D shape synthesis. Approaches without explicit structural modeling lack the ability to specify a particular desired topology, e.g. a chair with a particular slat configuration in the back. On the other hand, approaches that do model part-level structure are restricted to simple topologies defined by assemblies of a few coarse primitives and cannot model e.g. complex fretwork or ornamentation.
@@
Our approach implements the surface generation step by inferring and assembling a collection of locally-supported neural implicit functions, conditioned on the skeletal abstraction. We draw inspiration from recent work in this area which associates a latent code with each 3D point in a sparse set, and generates a local implicit from a latent grid (Zhang et al., 2022). The mixture of these implicits defines the overall synthesized shape, and allows for better generation of fine geometric details than a single large implicit. However, the sparse set of point supports in prior work tends to be arbitrary and not very interpretable. Follow-up work based on 3D neural fields and cross-attention (Zhang et al., 2023) drops explicit spatial grounding on the latent grids altogether. In contrast, our skeleton-based latent grids are more structure-aware, providing more interpretable supports for latent codes in 3D space, while remaining capable of representing complex, fine-grained topological structure.
@@ 2. Related work
Skeletonization
3D skeletonization, the computation of medial skeletons from a surface representation, is a well-studied problem in geometry processing. For a full discussion, we refer readers to the survey of Tagliasacchi et al. (2012). Skeletons, 1D curves, 2D sheets or a combination of them, have been used as an intermediate representation for creating complete shapes. Traditional optimization strategies have been proposed via property-grouping and sinking consolidation to extract skeletons (Tagliasacchi et al., 2009; Wu et al., 2015). More recently, Yin et al. (2018) proposed a neural network that can translate surface point clouds to skeletal point clouds, and Nie et al. (2020) followed up by regressing complete skeletons from partial surfaces. Clemot et al. (2023) fitted an implicit neural representation to surface points and extracted skeletons as the inner points with minimum SDF value that lie along the opposite direction to the gradient. On the other hand, our work is focused not only on computing those representations, but also on creating efficient shape generation models grounded on medial abstractions. While we also explore how to estimate such abstractions from point clouds, we primarily demonstrate how these intermediate representations can be used to guide the generation process, leading to more interpretable and topologically faithful shapes.
@@
3D shapes can be approximated as unions of spheres. For instance, a surface can be reconstructed by piecewise linear interpolation of medial spheres centered on skeleton points (Li et al., 2016). Shape modeling tools like ZBrush (Pixologic, 2022) allow artists to manually create such a skeleton of spheres via a feature called ZSpheres. Sphere Meshes (Thiery et al., 2013) used this representation for automatic shape approximation and extended it to non-tubular geometry. Convolution Surfaces (Bloomenthal and Shoemake, 1991) are another class of skeleton-based implicit surfaces where the surface is defined as the level set of a function obtained by integrating a kernel function along the skeleton. To avoid blending artifacts, Zanni et al. (2013) proposed scale-invariant blending. Our work investigates using such abstractions to guide the generative process for the best of two worlds – efficient topological representation and interpretability from medial abstractions, and detailed surface generation from data-driven neural fields.
We are inspired by Yin et al. (Yin et al., 2018) and Nie et al. (2020), who developed neural networks that translate (complete and partial, respectively) surface point clouds to meso-skeletons, and back to surface points from which continuous surfaces can be extracted with Poisson reconstruction. In contrast, we focus on the generation rather than the regression task, with consequent major differences in the technical design and applications. We also develop a novel neural surface representation based on skeleton-supported local implicits that directly yields a continuous surface, and is designed to be efficient, detailed, and artifact-free.
@@
Scalar fields parameterized by neural networks have been frequently used as a shape representation in many deep generative models (Mescheder et al., 2019; Park et al., 2019; Chen and Zhang, 2019). When compared to other representations like voxels or meshes, those neural fields are capable of representing shapes with varying topology with reasonable amount of detail without requiring prohibitive memory footprint. More recent approaches encode spatially-varying features that locally modulate the neural fields to improve the accuracy of the generated shapes (Zhang et al., 2022, 2023; Peng et al., 2020). Unfortunately, those approaches can generate topologically incorrect and hard to control shapes. Our approach addresses both of these issues by representing shapes as neural fields conditioned on medial abstractions. More specifically, we learn scalar fields that encode a “thickening” operation on a skeletal representation, that naturally lead to more interpretable, controllable and topologically-faithful shapes.
@@
A full review of 3D generative models is beyond our scope: we refer the reader to excellent surveys (Xu et al., 2023; Shi et al., 2023). Methods that directly generate “raw” representations such as voxels, meshes, or point clouds frequently yield topological artifacts, omissions, and other inaccuracies in regions with thin/fine details. Further, they lack interpretable intermediates to help users control the generative process. Structure-aware methods (Chaudhuri et al., 2020), in particular part-based models (e.g., (Li et al., 2017)), have potentially higher topological fidelity and interpretability. However, they are currently suitable for shapes with small numbers of meaningful parts, and not for more topologically complex structures with ambiguous part decompositions. We use medial abstractions as a structure-aware intermediate representation to avoid these limitations.
@@ 3. Method
Figure 3. A shape and its medial axis (purple) (a) medial balls (yellow) vs (b) enveloping primitives (Guo et al., 2023).
@@ 3.1. Preliminaries
Consider a closed, oriented, and bounded shape $S$ in $\mathbb{R}^{3}$, the medial axis is represented as a set of centers of maximal spheres inscribed within the shape (Figure 3, left). Each of these medial spheres is tangent to at least two points on the boundary $\partial S$ of S and does not contain any other boundary points in its interior. The medial axis transformation (MAT) comprises both the medial axis and the radius associated with each sphere center (Tagliasacchi et al., 2016). The MAT can be used to compactly capture both topological shape information and its geometry (Li et al., 2016), yet as also recently noted in (Guo et al., 2023), its ability to capture local surface details, such as shape protrusions and other high curvature regions, is limited, since a substantial number of medial spheres are often needed (Figure 3, left).
@@ 3.2. Surface generation
We
extend the above primitives to “neural envelopes”, such that can be used in our neural generative model. We employ a neural network, parameterized by learnable parameters $\theta$, to approximate the radius function. In our case, the radius function depends not only on the direction from a medial element to the query point but also a medial latent code, which is specific to the medial element and aims to encode surface information around the medial axis. The latent code helps decoding towards a more accurate surface, since it is trained to encode shape information. In addition, instead of finding the medial element closest to the query point, as done in (Guo et al., 2023), we found that a significantly better surface approximation can be achieved by finding the medial element whose surface envelope is closest to the query point (Figure 4, see also our ablation). Specifically, our neural enveloping function defines the following implicit:
@@
Given the surfaces of training shapes in a dataset along with their medial elements, the parameters of the MLP can be trained to minimize the $L_{1}$ loss measuring error between predicted locations of sample points and the location of training surface sample points sampled at different directions around each medial element:
@@ 3.3. Generative model
We developed a fast, distance field-based skeletonization algorithm whose basic idea is to move surface sample points set towards the negative gradient of the signed distance function through an iterative gradient descent procedure such that surface shrinks and moves towards the medial axis. We provide more details about our skeletonization in the supplementary material. We stress that the skeletonization was performed fully automatically for all our datasets without any manual parameter tuning per shape category or testing scenario.
@@
As an additional operation, each block adds a positional embedding to the input feature vector $\mathbf{h}_{i}$ of each medial point based on its position: $\mathbf{h}_{i}^{\prime(l)}=\mathbf{h}_{i}^{l}+g(\mathbf{s}_{i})$, where $g$ represents a frequency-based positional embedding of the medial point positions (Sitzmann et al., 2020), followed by a fully connected layer. We observed that adding this positional embedding offered significantly better reconstruction results, since it made each latent aware of its corresponding medial position represented at different frequencies. The feature representations of the last block is finally processed through a block of self-attention layers to exchange the information within the latent set – this last block produces the final latents used in our surface decoder (Eq. 2).
@@ 4. Results & Applications
We now discuss the experiments and validation of our method for three different applications: category-conditioned shape generation, surface reconstruction from point clouds, and skeleton-driven shape generation. We note that all our code, data, and evaluation will be publicly released.
@@ 4.1. Category-driven shape generation
In this application, the input is a given category (e.g., “lamp”, “chair”, and so on) and the output is a sample set of 3D generated shapes from this category. Our method here makes use of the diffusion stages discussed in Section 3.3: the first stage generates a point-based MAT given the input category and the second stage generates the medial latents. Then our surface decoder generates the surface under the guidance of the medial points and latents (Section 3.2).
Baselines
@@
In addition, as proposed in the literature of image generative models, we use the Fréchet Inception Distance and Kernel Inception Distance. While prior work (Zhang et al., 2023, 2022) relies on PointNet++ embeddings to evaluate these metrics, we use the 768-D embeddings taken by the more recent and more discriminative model of PointBert (Yu et al., 2022) trained on ShapeNetCore. We refer to these two metrics as PointBert-FID and PointBert-KID. Finally, we report generative precision and recall (Sajjadi et al., 2018) as alternative measures to assess how well generated data covers the reference test splits and vice versa. For precision and recall, the similarities are assessed via PointBert embeddings.
@@
We follow the experimental setup of 3DS2VS (Zhang et al., 2023), which was also demonstrated in the same application. We use the same dataset of ShapeNet-v2 (Chang et al., 2015) with the same splits. We use the same watertight ShapeNet meshes provided by (Zhang et al., 2022). In constrast to 3DS2VS that evaluated category-conditioned generation in $2$ categories, in our case we evaluate on the largest $10$ categories from ShapeNet, namely: tables, cars, chairs, airplanes, sofas, rifles, lamps, watercrafts, benches, and speakers. To further improve our evaluation, we generate a much larger set of generated samples – for each method and category, we generate a number of samples equal to $3$x the size of the ground truth test split per category, which is also consistent with what Achlioptas et al. (2018b) originally proposed.
@@
Metric 3DILG 3DS2VS GEM3D (Ours)
@@
Qualitative results
Figure 5 demonstrates generated samples from our method for various categories from ShapeNet. We observe that our method is able to generate structurally challenging patterns (e.g., grid-like patterns for chair backs), thin parts (e.g, wings for airplanes), tubular or generalized cylinder-like parts (e.g., lamp wires or other connecting pieces). Our teaser (Figure 1) shows additional results. For more sample visualizations, we refer to the supplement.
Timings
We note that our method takes approximately $30$ seconds to generate one shape measured on a NVidia RTX A6000.
4.2. Surface reconstruction from point clouds
In this application, the input to our method is a sparse point cloud of a shape and the output is a reconstructed surface. Here, we also follow the “auto-encoding” setting also used in 3DILG (Zhang et al., 2022) and 3DS2VS (Zhang et al., 2023), where the point cloud is obtained after sampling the original input surface. In this autoencoding setting, we only use the second diffusion stage of our method i.e., given a point-based skeleton, the second stage generates the medial latents, then our surface decoder uses them to reconstruct the surface. To apply our method on this task, we need as input a point-sampled skeleton. To obtain such a skeleton, we use a variant of P2PNet (Yin et al., 2018), an encoder-decoder network that takes as input a point cloud of a shape’s surface and converts it to a point-based skeleton representation. Compared to the original P2PNet, we replaced the original PointNet++ encoder (Qi et al., 2017) with the cross-attention-based shape encoder from 3DS2VS (Zhang et al., 2023), including its positional-based embedding functions. We train this P2PNet variant in the training split of ShapeNet. Then at test time, we use its output skeletons as input to our second diffusion stage. We emphasize that no manual input or supervision were needed to obtain the skeletons – P2PNet is trained in a unsupervised manner.
Baselines
We again compare our method with 3DILG (Zhang et al., 2022) and 3DS2VS (Zhang et al., 2023), which were also applied in the same auto-encoding setting. All methods use the same number of latent codes ($2048$ codes).
Metrics
For evaluation purposes, we compare reconstructed shapes to their corresponding reference (“ground-truth”) shapes, which the input point clouds were sampled from. Following 3DS2VS (Zhang et al., 2023), we use the (squared) Chamfer Distance (CD), Intersection Over Union (IoU) and F-Score (F1). Compared to 3DS2VS’ protocol, we made the following changes: (i) for computing IoU, we use a higher-resolution grid of $256^{3}$ instead of $128^{3}$ to better characterize reconstruction of small-scale surface details, thin parts, and holes, (ii) for CD and F1, we use farthest point sampling instead or random point sampling to better cover small and thing parts – random sampling tends to miss such parts. In the supplement, we provide comparisons using the original 3DS2VS’ protocol.
Experimental setup
Following (Zhang et al., 2022) and (Zhang et al., 2023), we use the same training and test split from ShapeNet-v2 to train and test our method respectively. Apart from ShapeNet, we also test all competing methods in a more challenging out-of-distribution testing scenario: after training on ShapeNet, we test all methods on Thingi-10K (Zhou and Jacobson, 2016). We note that the whole dataset is $10K$ test shapes, $4$ times larger than the ShapeNet’s test split. The Thingi10K shapes contain man-made objects that often possess highly complex topology (e.g., a large genus number). None of the Thingi10K shapes exists in ShapeNet. The vast majority of them do not even relate to any of the ShapeNet categories. Thingi10K also provides topological complexity information (genus) for a large subset of the dataset. We use this information as a proxy of shapes’ topological complexity for additional evaluation.
Quantitative results
Table 2 shows quantitative evaluation on ShapeNet-v2. The odd rows report the measures averaged over the $10$ largest categories of ShapeNet, and the even rows report the measures averaged over all $55$ categories. In terms of the surface-based metrics of CD and F1 scores, our model is consistently better than the baselines. In terms of IoU, our model has comparable performance. The results reflect a more accurate surface reconstruction, without compromising volumetric (IoU) accuracy. The supplement reports the performance for each of the largest $10$ categories – again, we outperform prior methods in most categories based on the surface-based metrics. The gap is particularly larger for topologically challenging categories, such as lamps and watercrafts.
The out-of-distribution evaluation of reconstruction on Thingi10K is shown in Table 3. Our model clearly outperforms the baselines – the performance gap increases for topologically challenging shapes with higher genus. We note that this evaluation does not include IoU since some Thingi10K ground-truth shapes are not watertight, thus, volume-based metrics cannot be accurately assessed.
Metric Test dataset 3DILG 3D2VS GEM3D (Ours)
CD ($\downarrow$) Shapenet (10 cat.) 1.69 1.32 1.17
CD ($\downarrow$) Shapenet (all cat.) 2.05 1.70 1.38
IOU ($\uparrow$) Shapenet (10 cat.) 90.1 96.2 95.9
IOU ($\uparrow$) Shapenet (all cat.) 90.5 95.4 95.5
F1 ($\uparrow$) Shapenet (10 cat.) 97.7 98.5 99.3
F1 ($\uparrow$) Shapenet (all cat.) 96.3 97.4 98.7
Table 2. Evaluation measures on point cloud surface reconstruction (“auto-encoding”) in ShapeNet. The rows “ShapeNet (10 cat.)” report the measures averaged over the $10$ largest categories of ShapeNet, while the rows “ShapeNet (all)” report the measures averaged over all $55$ ShapeNet categories. For evaluation per-category, please see the supplement.
Thingi10K #$\#shapes$ 3DILG 3D2VS GEM3D (Ours)
all 9997 2.12 1.74 1.68
$genus\geq 0$ 5608 1.93 1.62 1.61
$genus\geq 5$ 1037 2.41 1.96 1.78
$genus\geq 10$ 489 2.88 2.30 1.98
CD $\downarrow$ $genus\geq 20$ 229 3.67 2.91 2.32
all 9997 93.6 96.4 96.7
$genus\geq 0$ 5608 95.3 97.4 97.4
$genus\geq 5$ 1037 91.1 94.8 95.9
$genus\geq 10$ 489 87.5 92.3 93.7
F1 $\uparrow$ $genus\geq 20$ 229 82.6 88.0 90.4
Table 3. Evaluation measures on point cloud surface reconstruction in the Thingi10K dataset. We note that none of the methods are trained on Thingi10K, or any subset of it. We report performance in terms of CD and F1 scores for shapes of increasing genus.
Qualitative results
Figures 6 and 7 shows qualitative evaluation on ShapeNet-v2 and Thingi10K respectively. In particular, the Thingi10K results show that our method can better reconstruct surface topology e.g., the connectivity of different shape parts and their surface holes. We suspect that this is due to the ability of our method to capture structure and topology information through the neural medial representation, which in turn guides the reconstruction. In contrast, 3DILG and 3D2VS tend to oversmooth surface details, miss connections, and alter the topology in an unpredicted manner.
4.3. Skeleton-driven shape synthesis
An alternative application of our method is to generate surfaces driven by an input skeleton in the form of a point-based MAT. From a practical point of view, one possibility for users of our method is to execute our first diffusion stage, obtain various shape structures represented in their MATs, then select the one matching their target structure or topology goals more. Then the user can execute the second stage to obtain various shape alternatives conforming to the given structure. Figure 8 highlights several examples of this application scenario. We show a generated MAT, then diverse surfaces conforming to this MAT. For example, our method can generate diverse lamp stands from a table lamp structure or chairs of diverse part thicknesses conforming to a particular chair structure.
4.4. Ablation studies
In our supplementary material, we show the impact of various choices in our method, including testing with different number of medial points and using closest medial points versus closest envolopes in our implicit function.

### 11 Apr 2024 v2

@@ Abstract.
We introduce GEM3D  [^1] Project page (with code): lodurality.github.io/GEM3D – a new deep, topology-aware generative model of 3D shapes. The key ingredient of our method is a neural skeleton-based representation encoding information on both shape topology and geometry. Through a denoising diffusion probabilistic model, our method first generates skeleton-based representations following the Medial Axis Transform (MAT), then generates surfaces through a skeleton-driven neural implicit formulation. The neural implicit takes into account the topological and geometric information stored in the generated skeleton representations to yield surfaces that are more topologically and geometrically accurate compared to previous neural field formulations. We discuss applications of our method in shape synthesis and point cloud reconstruction tasks, and evaluate our method both qualitatively and quantitatively. We demonstrate significantly more faithful surface reconstruction and diverse shape generation results compared to the state-of-the-art, also involving challenging scenarios of reconstructing and synthesizing structurally complex, high-genus shape surfaces from Thingi10K and ShapeNet.
@@
Figure 1. We present GEM3D, a neural generative model able to (a) synthesize skeleton-based representations, or in other words, medial abstractions (in red), and surfaces (in green) that follow the structure and topology information encoded in the skeletons. GEM3D can also be used in various applications, such as (b) reconstructing topologically complex surfaces, (c) generating diverse shape samples following a given skeleton, and (d) generating surfaces from novel, user-specified skeletons.
@@ 1. Introduction
Automated or semi-automated 3D shape synthesis is a significant and challenging problem in geometric modeling, with wide-ranging applications to computer-aided design, fabrication, architecture, art, and entertainment. While early work in this space primarily focused on handwritten models (Prusinkiewicz and Lindenmayer, 1990; Musialski et al., 2013; Merrell et al., 2010), subsequent work employed statistical learning to infer generative design principles from data (Kalogerakis et al., 2012; Fisher et al., 2011). In recent years, a variety of approaches have developed deep neural network-based architectures for 3D synthesis (Xu et al., 2023; Shi et al., 2023; Chaudhuri et al., 2020; Patil et al., 2020). While these methods can capture diverse macro-level appearances, they rarely model shape structure or topology explicitly, relying instead on the representational power of the network to generate plausible-looking voxel grids (Liu et al., 2017), point clouds (Achlioptas et al., 2018a), meshes (Dai and Nießner, 2019), or implicit fields (Chen and Zhang, 2019). Since 3D networks are hampered by the additional resource overheads incurred by the extra dimension compared to 2D image generation networks, they often struggle to model fine detail and connectivity. Some approaches model part layouts (Li et al., 2017), but are limited in the complexity of the structures they can generate.
At the same time, these prior 3D synthesis methods rarely give artists flexible, precise control. They act more as black boxes for unconditional generation, or reconstruction from images or 3D scans. Recent methods introduce synthesis based on text prompts (Poole et al., 2023; Lin et al., 2023), with remarkable results but only high-level, global control via prompt engineering. 3D character artists have long been accustomed to posing skeleton rigs for accurate character configuration. However, such direct local control and interpretability through intuitive abstractions has had limited success in general 3D shape synthesis. Approaches without explicit structural modeling lack the ability to specify a particular desired topology, e.g. a chair with a particular slat configuration in the back. On the other hand, approaches that do model part-level structure are restricted to simple topologies defined by a few coarse primitives and cannot model complex fretwork or ornamentation.
@@
Our approach implements the surface generation step by inferring and assembling a collection of locally-supported neural implicit functions, conditioned on the skeletal abstraction. We draw inspiration from recent work in this area which associates a latent code with each 3D point in a sparse set, and generates a local implicit from a latent grid (Zhang et al., 2022). The mixture of these implicits defines the overall synthesized shape, and allows for better generation of fine geometric details than a single large implicit. However, the sparse set of point supports in prior work tends to be arbitrary and not very interpretable. Follow-up work based on 3D neural fields and cross-attention (Zhang et al., 2023) drops explicit spatial grounding on the latent grids altogether. In contrast, our skeleton-based latent grids are more structure-aware, providing interpretable supports for latent codes in 3D space, while remaining capable of representing complex, fine-grained topological structure.
@@ 2. Related work
3D Skeletonization
The computation of medial skeletons from a surface representation, is a well-studied problem in geometry processing. 1D curves or 2D sheets or a combination of them have been traditionally used as skeletal representations of 3D shapes. Most skeletonization methods rely on analytic approximations to skeletons, such as Voronoi diagrams (Amenta and Bern, 1998; Yan et al., 2018) and Power diagrams (Wang et al., 2022), morphological operations, such as surface contraction (Tagliasacchi et al., 2012), topological graphs, such as Reeb graphs (Hilaga et al., 2001), or optimization based on various geometric criteria (e.g., symmetry preservation (Tagliasacchi et al., 2009), normal preservation (Wu et al., 2015), set coverage (Dou et al., 2022)). For a full discussion of geometric methods for skeletonization, we refer readers to the survey by Tagliasacchi et al. (2016). More recently, neural network approaches have been proposed for extracting skeleton-based representations from point clouds  (Yin et al., 2018; Ge et al., 2023), implicit shape representations (Rebain et al., 2021; Clémot and Digne, 2023), volumetric representations (Xu et al., 2019), meshes (Xu et al., 2020), point cloud sequences (Xu et al., 2022a), or images (Hu et al., 2023, 2022; Xu et al., 2022b). Our work is focused not on computing those representations, but on creating efficient shape generation models grounded on medial abstractions. While we also explore how to estimate such abstractions from point clouds, we primarily demonstrate how these intermediate representations can be used to guide the generation process, leading to more interpretable and topologically faithful shapes.
@@
3D shapes can be approximated as unions of spheres. For instance, a surface can be reconstructed by piecewise linear interpolation of medial spheres centered on skeleton points (Li et al., 2016). Shape modeling tools like ZBrush (Pixologic, 2022) allow artists to manually create such a skeleton of spheres via a feature called ZSpheres. Sphere Meshes (Thiery et al., 2013) used this representation for automatic shape approximation and extended it to non-tubular geometry. Convolution Surfaces (Bloomenthal and Shoemake, 1991) are another class of skeleton-based implicit surfaces where the surface is defined as the level set of a function obtained by integrating a kernel function along the skeleton. To avoid blending artifacts, Zanni et al. (2013) proposed scale-invariant blending. Our work investigates using such abstractions to guide the generative process for the best of two worlds – efficient topological representation and interpretability from medial abstractions, and detailed surface generation from data-driven neural fields. More recently, neural networks have been developed to translate complete or partial surface point clouds (Yin et al., 2018; Nie et al., 2020; Yang et al., 2020) or RGB images (Tang et al., 2019, 2021; Hu et al., 2022) to skeletal representations, and back to either surface points or 3D meshes. None of these approaches are generative i.e., are capable of generating multiple valid skeletal representations given their inputs. We also develop a novel neural surface representation based on skeleton-supported local implicits that directly yields a continuous surface, and is designed to be efficient, detailed, and artifact-free.
@@
Scalar fields parameterized by neural networks have been frequently used as a shape representation in many deep generative models (Mescheder et al., 2019; Park et al., 2019; Chen and Zhang, 2019). When compared to other representations like voxels or meshes, those neural fields are capable of representing shapes with varying topology with reasonable amount of detail without requiring prohibitive memory footprint. More recent approaches encode spatially-varying features that locally modulate the neural fields to improve the accuracy of the generated shapes (Zhang et al., 2022, 2023; Peng et al., 2020; Hui et al., 2022; Cheng et al., 2023; Shue et al., 2023; Zheng et al., 2023; Zhou et al., 2023). These approaches can generate topologically incorrect shapes. Our approach addresses these issues by representing shapes as neural fields that encode a “thickening” operation on a medial skeletal representation, that naturally leads to more interpretable and topologically-faithful shapes.
@@
A full review of 3D generative models is beyond our scope: we refer the reader to excellent surveys (Xu et al., 2023; Shi et al., 2023). Methods that directly generate “raw” representations such as voxels, meshes, or point clouds frequently yield topological artifacts, omissions, and other inaccuracies in regions with thin/fine details. Further, they lack interpretable intermediates to help users control the generative process. Structure-aware representations (Chaudhuri et al., 2020), such as part-based models (e.g., (Huang et al., 2015; Li et al., 2017; Dubrovina et al., 2019; Wu et al., 2020; Petrov et al., 2023)) potentially have better topological fidelity and interpretability. However, they are suitable for shapes with small numbers of parts and not for more topologically complex structures with ambiguous part decompositions. We use medial abstractions as structure-aware intermediate representation to avoid these limitations.
@@ 3.1. Preliminaries
Consider a closed, oriented, and bounded shape $S$ in $\mathbb{R}^{3}$, the medial axis is represented as a set of centers of maximal spheres inscribed within the shape (Figure 3, left).
Figure 3. A shape and its medial axis (purple) (a) medial balls (yellow) vs (b) enveloping primitives.
Each of these medial spheres is tangent to at least two points on the boundary $\partial S$ of S and does not contain any other boundary points in its interior. The medial axis transformation (MAT) comprises both the medial axis and the radius associated with each sphere center (2016). The MAT can be used to compactly capture both topological shape information and its geometry (Li et al., 2016), yet as recently noted in (Guo et al., 2023), its ability to capture local surface details, such as shape protrusions and high curvature regions, is limited, since a substantial number of medial spheres are often needed (Figure 3, left).
@@ 3.2. Surface generation
We extend the above primitives to “neural envelopes”, such that can be used in our neural generative model. We employ a neural network, parameterized by learnable parameters $\theta$, to approximate the radius function. In our case, the radius function depends not only on the direction from a medial element to the query point but also a medial latent code, which is specific to the medial element and aims to encode surface information around the medial axis. The latent code helps decoding towards a more accurate surface, since it is trained to encode shape information. In addition, instead of finding the medial element closest to the query point, as done in (Guo et al., 2023), we found that a significantly better surface approximation can be achieved by finding the medial element whose surface envelope is closest to the query point (Figure 4, see also our ablation). Specifically, our neural enveloping function defines the following implicit:
@@
Given the surfaces of training shapes with their medial elements, the parameters of the MLP can be trained to minimize the $L_{1}$ loss measuring error between predicted locations of sample points and the location of training surface sample points sampled at different directions around each medial element:
@@ 3.3. Generative model
We developed a fast, distance field-based skeletonization algorithm whose basic idea is to move surface sample points set towards the negative gradient of the signed distance function through an iterative gradient descent procedure such that surface shrinks and moves towards the medial axis. We provide more details about our skeletonization in the supplementary material. The skeletonization was performed automatically without any manual parameter tuning per shape category or testing scenario.
@@
As an additional operation, each block adds a positional embedding to the input feature vector $\mathbf{h}_{i}$ of each medial point based on its position: $\mathbf{h}_{i}^{\prime(l)}=\mathbf{h}_{i}^{l}+g(\mathbf{s}_{i})$, where $g$ represents a frequency-based positional embedding of the medial point positions (Sitzmann et al., 2020), followed by a fully connected layer. We observed that adding this positional embedding offered significantly better reconstruction results, since it made each latent aware of its corresponding medial position. The feature representations of the last block is processed through a block of self-attention layers to exchange the information within the latent set – it produces the final latents used in our surface decoder (Eq. 2).
@@ 4. Results & Applications
We discuss the experiments and validation of our method for three applications: category-conditioned shape generation, surface reconstruction from point clouds, and skeleton-driven shape generation.
@@ 4.1. Category-driven shape generation
In this application, the input is a given category (e.g., “lamp”, “chair”, and so on) and the output is a sample set of 3D generated shapes from this category. Our method here makes use of the diffusion stages discussed in Section 3.3: the first stage generates a point-based MAT given the input category and the second stage generates the medial latents given the input category. Then our surface decoder generates the surface under the guidance of the medial points and latents (Section 3.2).
Baselines and metrics
@@
In addition, as proposed in the literature of image generative models, we use the Fréchet Inception Distance and Kernel Inception Distance. While prior work (Zhang et al., 2023, 2022) relies on PointNet++ embeddings to evaluate these metrics, we use the 768-D embeddings taken by the more recent and more discriminative model of PointBert (Yu et al., 2022). We refer to these two metrics as PointBert-FID and PointBert-KID. Finally, we report generative precision and recall (Sajjadi et al., 2018) as alternative measures to assess how well generated data covers the reference test splits and vice versa. For precision and recall, the similarities are assessed via PointBert embeddings.
@@
We follow the experimental setup of 3DS2VS (Zhang et al., 2023), which was also demonstrated in the same application. We use the same dataset of ShapeNet-v2 (Chang et al., 2015) with the same splits. We use the same watertight ShapeNet meshes provided by (Zhang et al., 2022). In constrast to 3DS2VS that evaluated category-conditioned generation in $2$ categories, in our case we evaluate on the largest $10$ categories from ShapeNet, namely: tables, cars, chairs, airplanes, sofas, rifles, lamps, watercrafts, benches, and speakers. To further improve our evaluation, for each method and category we generate a number of samples equal to $3$x the size of the ground truth test split per category, which is also consistent with Achlioptas et al. (2018b) originally proposed.
@@
Metric 3DILG 3DS2VS Ours